I have a problem where the perimeter and area of a rectangle are the same. The width of the rectangle is 3 and the length is x. I need to write an equation to solve for x to get the area and perimeter. I have tried but I keep coming up with equations that have two varibles, instead of just one (x), and I haven't yet learned how to find the value of x if there is another variable in the problem instead of two other numbers.
Answer by ankor@dixie-net.com(22740)

a problem where the perimeter and area of a rectangle are the same. The width of the rectangle is 3 and the length is x. I need to write an equation to solve for x to get the area and perimeter.
:
Area = perimeter
L*W = 2L + 2W
Given that W=3
3L = 2L + 2(3)
3L - 2L = 6
L = 6 is the length
See if that checks out
3*6 = 2(6) + 2(3)
18 = 12 + 6